What is the length of a rectangular prism when volume is 126, width is 7 7/8, & height is 2?
Answer by Alan3354(69443)

What is the length of a rectangular prism when volume is 126, width is 7 7/8, & height is 2?
======================
Vol = L*W*H = 126
L*7.875*2 = 126
L = 126/15.75